Year/speed 2008/1028 2009/1727 2010/2208 2011/2704 A) find the equation of your trendline using ...

year/speed
2008/1028
2009/1727
2010/2208
2011/2704

A) find the equation of your trendline using the mean point an one other point that is on your line

b)find the equation for the best fit line using a linear regression tool.
Answer #1

y/ OD 1028 172:7 201, 12-70% Mean 2009.5| 1916.75 Me ano = 8038/年=오009.5 Х 2.32-25 133-12s 94-875 (4562一 (80-87 787-25 2 -2S2019 4-2 14:151727 2010 2206 2011 Mean X = 2009.5 Meany= 1916-75- 1.5 -0.5 0.5 I.5 828.75 -18975 291.25 787,25 1333. 125 44.375 195 625 11